# Computing Richelot isogeny graph of superspecial abelian threefolds

### v1

@@ 2.4. Superspecial abelian threefold isogeny graph
We note that if two curves $C,C^{\prime}$ are not isomorphic to each other, then neither are their Jacobians (as principally polarized abelian varieties) from Torelli theorem. Let $\Lambda_{g,p}$ be the list of isomorphism classes of superspecial genus-$g$ curves over $\overline{\mathbb{F}}_{p}$. For $g\leq 3$, the number #$\#\Lambda_{g,p}$ is given by [11], [21], and [3], however it is not known how many of those curves are hyperelliptic, especially. In the following, we explain how to generate the list $\Lambda_{g,p}$ by using the isogeny graph. First of all, we define the isogeny graphs of superspecial abelian varieties:
@@
Let $\ell$ be a prime integer. The superspecial isogeny graph $\mathcal{G}_{g}(\ell,p)$ is the directed multigraph obtained as
@@
In the following, we discuss the case where $\ell=2$, then the graph $\mathcal{G}_{g}(2,p)$ is often called the superspecial Richelot isogeny graph. The list $\Lambda_{g,p}$ can be generated using the superspecial Richelot isogeny graph as in Algorithm 2.1:
1:An integer $g\geq 2$ and a prime integer $p\geq 3$
2:The list of isomorphism classes of all superspecial curves of genus $g$
3:Generate the list $\Lambda_{1,p}$ of isom. classes of s.sing. elliptic curves
4:Let $\mathcal{L}$ be the list of isom. classes of $E_{1}\times\cdots\times E_{g}$ for all $E_{i}\in\Lambda_{1,p}$ and $k\leftarrow 1$
5:while $k\leq\#\mathcal{L}$ do
6:     Let $A$ be the $k$-th element of $\mathcal{L}$
7:     Compute all codomains of Richelot isogenies whose domain is $A$
8:     for all codomains $A^{\prime}$ obtained in the previous step do
9:         if $A^{\prime}$ is not isomorphic to any element of $\mathcal{L}$ then
10:              Add $A^{\prime}$ to the end of $\mathcal{L}$
11:         end if
12:     end for
13:     Let $k\leftarrow k+1$
14:end while
15:return the list of genus-$g$ curves $C$ such that the Jacobian of $C$ belongs to $\mathcal{L}$
Algorithm 2.1 ListUpSuperspecialCurves
The correctness of this algorithm can be checked thanks to the following theorem:
Theorem 2.20 ([24, Section 7]).
@@
In the case of $g=2$, it is easy to implement Algorithm 2.1 (indeed, this is used in [25, Algorithm 5.1]). We can compute Richelot isogenies in Step 5 using formulas in [19, Proposition 4] and [38, Chapter 8]. Moreover, isomorphism tests in Step 7 can be done by computing Igusa invariants of a genus-2 curve (cf. [22]).
@@
The structure of the graph $\mathcal{G}_{2}(2,p)$ was studied by many authors (cf. [27], [15] and so on). An important fact is that the Jacobians’ subgraph $\mathcal{J}_{2}(2,p)$ of $\mathcal{G}_{2}(2,p)$ is also connected by [14, Theorem 7.2]. Here $\mathcal{J}_{g}(\ell,p)$ is defined to be
  • The vertices are isomorphism classes of all the Jacobians of genus-$g$ curves over $\overline{\mathbb{F}}_{p}$.
@@
These researches may make it possible for us to construct a more efficient algorithm for listing up all superspecial genus-2 curves rather than Algorithm 2.1.
Similarly to the case of $g=2$, we want to list up all genus-$g$ superspecial curves for $g\geq 3$. However, no explicit formula is known for the computations of $(2,\ldots,2)$-isogenies in general unlike the genus-2 case. In the next section, we will propose an algorithm for computing $(2,2,2)$-isogenies using theta functions.
@@ 3. Computing the Richelot isogeny graph of dimension $3$
In this section, we give an explicit algorithm (Algorithm 3.7) for computing the graph $\mathcal{G}_{3}(2,p)$. Our algorithm uses theta functions introduced in Section 2.2, but is still valid over a field of odd characteristic $p\geq 3$, thanks to the algebraic theory by Mumford [31]. The following three subsections describe the subroutines used in our main algorithm as follows:
  • Section 3.1: For a given theta null-point of an abelian threefold $A$, compute a theta null-point of one $A^{\prime}$, which is $(2,2,2)$-isogenous to $A$.
@@
At the last of this section (Section 3.4), we will combine them and write down our main algorithm explicitly.
@@ 3.1. How to compute one path from the vertex
In this subsection, let $A$ be a complex abelian threefold and $\varOmega\in\mathcal{H}_{3}$ be a period matrix of $A$. First of all, we prepare the following fact about a theta null-point of $A$.
@@
For all matrices $\varOmega\in\mathcal{H}_{3}$, we have the equation
$\vartheta_{0}\vartheta_{1}\vartheta_{2}\vartheta_{3}-\vartheta_{4}\vartheta_{5}\vartheta_{6}\vartheta_{7}=\vartheta_{32}\vartheta_{33}\vartheta_{34}\vartheta_{35}$
with the notation in Definition 2.11.
@@
$m_{3}$ $\coloneqq\begin{pmatrix}0&0&0\\ 0&1/2&0\end{pmatrix},\quad$ $m_{4}$ $\coloneqq\begin{pmatrix}0&0&0\\ 1/2&1/2&0\end{pmatrix},\vspace{0.5mm}$
@@
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(1,1,1,0,0,0)$ in the case of $k=1$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,1,0,1,1,0)$ in the case of $k=2$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,1,0,0,0,0)$ in the case of $k=3$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,1,0,1,1)$ in the case of $k=4$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,1,0,0,0)$ in the case of $k=5$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,0,0,1,0)$ in the case of $k=6$
@@
$\alpha\coloneqq\delta\coloneqq\mathbf{1}_{3},\quad\beta\coloneqq\mathbf{0}_{3},\quad\gamma\coloneqq\begin{pmatrix}4c_{11}&2c_{12}&2c_{13}\\ 2c_{12}&4c_{22}&2c_{23}\\ 2c_{13}&2c_{23}&4c_{33}\end{pmatrix}.\vspace{0.5mm}$
Then, one can check that $M$ belongs to ${\rm Sp}_{6}(\mathbb{Z})$ by Lemma 2.2, and moreover $M$ is the matrix we desire since
@@
(3.6) $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,2\varOmega)^{2}=\frac{1}{2^{3}}\hskip 0.56905pt\sum_{\beta\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)\theta\bigl{[}\begin{smallmatrix}0\\ b+\beta\end{smallmatrix}\bigr{]}(0,\varOmega)\vspace{0.5mm}$
for all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}$. Since the right-hand side of (3.6) is obtained by the values $\vartheta_{0},\ldots,\vartheta_{7}$, we can compute a theta null-point of $B$. As a summary of the statements in this subsection, we give an explicit algorithm (Algorithm 3.2) to compute a null-point of the codomain of a $(2,2,2)$-isogeny:
@@
It suffices to do the loops from Step 8 to Step 14 only for $i\in I_{\rm even}$ since we have that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$ for all $i\in I_{\rm odd}$, in fact.
@@ 3.2. How to compute all paths from the vertex
$\bigl{\{}M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{6}(\mathbb{Z})\mid\gamma\equiv\mathbf{0}_{3}\!\pmod{n}\bigr{\}}\vspace{0.5mm}$
@@
Remark 3.5.
There exist exactly $N_{3}(2)=135$ different $(2,2,2)$-isogenies from $A$ by Proposition 2.4, and thus the order of ${\rm Sp}_{6}(\mathbb{Z})/\varGamma_{0}(2)$ is equal to $135$.
@@
In the following, let $\mathcal{M}$ be the output of the function CollectSymplecticMatrices. Then, we can compute theta null-points of $135$ abelian threefolds $(2,2,2)$-isogenous to $A$ by using Theorem 2.9.
@@ 3.3. How to restore a defining equation of a genus-3 curve
  • If $N_{\rm van}=1$ and $S_{\rm van}\neq\{61\}$, how should we restore a definind equation of the corresponding genus-3 curve?
@@
$P_{i}\coloneqq\begin{pmatrix}\mathbf{1}_{3}&\beta_{i}\\[-0.85358pt] \gamma_{i}&\mathbf{1}_{3}\end{pmatrix},\vspace{-1mm}$
@@
    $\beta_{i}\coloneqq{\rm diag}(b_{1},b_{2},b_{3}),\quad\gamma_{i}\coloneqq{\rm diag}(a_{1},a_{2},a_{3}).\vspace{0.5mm}$
@@
    $\beta_{27}\coloneqq{\small\begin{pmatrix}1&1&0\\ 1&1&0\\ 0&0&0\end{pmatrix}},\quad\gamma_{27}\coloneqq{\small\begin{pmatrix}1&-1&0\\ -1&1&0\\ 0&0&0\end{pmatrix}}.$
@@
    $\beta_{62}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix}},\quad\gamma_{62}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}}.$
@@
In the following, let $A$ be an abelian threefold isomorphic to the Jacobian of a genus-3 hyperelliptic curve $C$, and we assume that a theta null-point of $A$ is given. In this setting, there uniquely exists $i\in I_{\rm even}$ such that $\vartheta_{i}=0$ by Proposition 2.15. If $i=61$, we can apply Proposition 2.16 for restoring a defining equation of $C$, but otherwise not. Here, let $M:=P_{61}{P_{i}}^{-1}\in{\rm Sp}_{6}(\mathbb{Z})$ and $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ be the theta null-point after acting $M$ on given $[\vartheta_{i}^{2}]_{i}$, and we have ${\vartheta^{\prime}}_{\!\!61}=0$ by Lemma 3.8. Then we can apply Proposition 2.16 to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$, which enables us to recover $C$ (Algorthm 3.5).
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=1$ of an abelian threefold $A$
@@
[…]
@@
By using these equations, we can restore a defining equation of a plane quartic $C$, as in Algorithm 3.6.
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=0$ of an abelian threefold $A$
@@
3.4. Algorithm for computing the graph $\mathcal{G}_{3}(2,p)$
In this subsection, let $p$ be a prime integer greater than $7$ and we give an explicit algorithm for computing the superspecial Richelot isogeny graph $\mathcal{G}_{3}(2,p)$. First, we need to prepare the list $\Lambda_{2,p}$ of all isomorphism classes of superspecial genus-2 curves, in addition to the list $\Lambda_{1,p}$ of all isomorphism classes of supersingular elliptic curves. Using these lists, we can compute the graph $\mathcal{G}_{3}(2,p)$ as in Algorithm 3.7.
1:A prime integer $p\geq 11$ and the lists $\Lambda_{1,p}$ and $\Lambda_{2,p}$
2:The lists representing all vertices and all nodes of the graph $\mathcal{G}_{3}(2,p)$
3:Initialize $\mathcal{V}_{1}\leftarrow\emptyset,\,\mathcal{V}_{2}\leftarrow\emptyset,\,\mathcal{V}_{3}\leftarrow\emptyset,\,\mathcal{V}_{4}\leftarrow\emptyset$ and $\mathcal{S}\leftarrow\emptyset$
4:for all $E_{1},E_{2},E_{3}\in\Lambda_{1,p}$ do
5:     Let $j_{1},j_{2},j_{3}$ be the $j$-invariant of $E_{1},E_{2},E_{3}$
6:     Compute a theta null-point $[\vartheta^{2}_{i}]_{i}$ of the abelian threefold $E_{1}\times E_{2}\times E_{3}$
7:     Add $inv\coloneqq(j_{1},j_{2},j_{3})$ to the list $\mathcal{V}_{4}$ and $\bigl{(}[\vartheta^{2}_{i}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the list $\mathcal{S}$
8:end for
9:for all $E\in\Lambda_{1,p},\hskip 0.85358ptC\in\Lambda_{2,p}$ do
10:     Let $j$ be the $j$-invariant of $E$, and $j^{\prime}$ be the Igusa invariant of $C$
11:     Compute a theta null-point $[\vartheta^{2}_{i}]_{i}$ of the abelian threefold $E\times{\rm Jac}(C)$
12:     Add $inv\coloneqq(j,j^{\prime})$ to the list $\mathcal{V}_{3}$ and $\bigl{(}[\vartheta^{2}_{i}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the list $\mathcal{S}$
13:end for
14:Initialize $\mathcal{E}\leftarrow\emptyset$ and let $k\leftarrow 1$
15:repeat
16:     Let $\bigl{(}[\vartheta^{2}_{i}]_{i},inv\bigr{)}$ be the $k$-th element of $\mathcal{S}$
17:     for all $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}\in\hskip 2.27621pt$ComputeAllRichelotIsogenies($[\vartheta^{2}_{i}]_{i}$do
18:         Compute the number $N_{\rm van}$ of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$
19:         if $N_{\rm van}=0$ then
20:              Let $inv^{\prime}$ be the Dixmier-Ohno invariant of ${\sf RestorePlaneQuartic}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
21:              if $inv^{\prime}\notin\mathcal{V}_{1}$ then
22:                  Add $inv^{\prime}$ to the list $\mathcal{V}_{1}$ and $\bigl{(}[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the end of $\mathcal{S}$
23:              end if
24:         else if $N_{\rm van}=1$ then
25:              Let $inv^{\prime}$ be the Shioda invariant of ${\sf RestoreHyperellipticCurve}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
26:              if $inv^{\prime}\notin\mathcal{V}_{2}$ then
27:                  Add $inv^{\prime}$ to the list $\mathcal{V}_{2}$ and $\bigl{(}[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the end of $\mathcal{S}$
28:              end if
29:         else if $N_{\rm van}=6$ then
30:              Restore $E\in\Lambda_{1,p}$ and $C\in\Lambda_{2,p}$, s.t. $E\times{\rm Jac}(C)$ corresponds to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$
31:              Let $inv^{\prime}:=(j,j^{\prime})$ where $j(E)=j$ and the Igusa invariant of $C$ is $j^{\prime}$
32:         else
33:              Restore $E_{1},E_{2},E_{3}\in\Lambda_{1,p}$, s.t. $E_{1}\times E_{2}\times E_{3}$ corresponds to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$
34:              Let $inv^{\prime}:=\{j_{1},j_{2},j_{3}\}$ where $j(E_{1})=j_{1},\,j(E_{2})=j_{2}$ and $j(E_{3})=j_{3}$
35:         end if
36:         Add $(inv,inv^{\prime})$ to the list $\mathcal{E}$
37:     end for
38:     Set $k\leftarrow k+1$
39:until $k\hskip 0.56905pt>\#\mathcal{S}$
40:return $\mathcal{V}_{1}\sqcup\mathcal{V}_{2}\sqcup\mathcal{V}_{3}\sqcup\mathcal{V}_{4}$ and $\mathcal{E}$
Algorithm 3.7  ComputeSspRichelotIsogenyGraph
Recall from Section 2.3 that an abelian threefold is classified as Type $1,2,3$ or $4$ and $\mathcal{V}_{i}$ in the above algorithm represent the list of vertices of each type, respctively. In Steps $4$ and $9$, we have to compute a theta null-point of an abelian threefold of Type $3$ and $4$. Here, in the following, we explain how to obtain them briefly:
Lemma 3.9.
Let $E=\mathbb{C}/(\mathbb{Z}+\tau\mathbb{Z})$ with $\tau\in\mathbb{C}$ be an elliptic curve over $\mathbb{C}$ defined by the equation $y^{2}=x(x-1)(x-t)$. Then, there exists $c\in\mathbb{C}^{\times}$ such that
$\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c,\quad\theta\bigl{[}\begin{smallmatrix}0\\ 1/2\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c\sqrt{t},\quad\theta\bigl{[}\begin{smallmatrix}1/2\\ 0\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c\sqrt{1-t}$
and moreover $\theta\bigl{[}\begin{smallmatrix}1/2\\ 1/2\end{smallmatrix}\bigr{]}(0,\tau)=0$.

### v2

@@ 2.4. Superspecial Richelot isogeny graph
We note that if two curves $C,C^{\prime}$ are not isomorphic to each other, then neither are their Jacobians (as principally polarized abelian varieties), by Torelli’s theorem. Let $\Lambda_{g,p}$ be the list of isomorphism classes (over $\overline{\mathbb{F}}_{p}$) of superspecial genus-$g$ curves. The numbers #$\#\Lambda_{g,p}$ for $g\leq 3$ are given by [12], [24], and [21], respectively (see [30] for a survey). However, it is not known how many of those curves are hyperelliptic for genus $g=3$. In the following, we explain how to generate the list $\Lambda_{g,p}$, by using the isogeny graph of superspecial abelian varieties:
@@
Let $\ell$ be a prime integer. The superspecial isogeny graph $\mathcal{G}_{g}(\ell,p)$ is the directed multigraph with the following properties
@@
The fact that the Pizer’s graph $\mathcal{G}_{1}(\ell,p)$ is connected is well-known (cf. [44]), but more surprisingly, the same assertion holds for $\mathcal{G}_{g}(\ell,p)$ as a generalization:
Theorem 2.20.
@@
Proof.
See [27, Theorem 34] or [1, Corollary 2.8]. ∎
In the following, we discuss the case where $\ell=2$, then the graph $\mathcal{G}_{g}(2,p)$ is often called the superspecial Richelot isogeny graph. Then, the list $\Lambda_{2,p}$ can be generated by using $\mathcal{G}_{2}(2,p)$ as in the following pseudocode (cf. [31, Algorithm 5.1]).
1:A prime integer $p>5$ and the list $\Lambda_{1,p}$ of isom. classes of s.sing. ell. curves
2:The list $\Lambda_{2,p}$ of isom. classes of s.sp. genus-2 curves
3:Let $\mathcal{L}$ be the list of isom. classes of $E_{1}\times E_{2}$ with $E_{1},E_{2}\in\Lambda_{1,p}$, and set $k\leftarrow 1$
4:repeat
5:     Let $A$ be the $k$-th abelian surface of $\mathcal{L}$
6:     for all abelian surfaces $A^{\prime}$ which are $(2,2)$-isogenous to $A$ do
7:         if $A^{\prime}$ is not isomorphic to any abelian surface of $\mathcal{L}$ then
8:              Add $A^{\prime}$ to the end of $\mathcal{L}$
9:         end if
10:     end for
11:     Set $k\leftarrow k+1$
12:until $k>\#\mathcal{L}$
13:return  the list of genus-2 curves whose Jacobian belongs to $\mathcal{L}$
Algorithm 2.21 ListUpSspGenus2Curves
In Algorithm 2.21, one can compute the codomains of Richelot isogenies of line 4 using formulae in [22, Proposition 4] and [49, Chapter 8]. In addition, isomorphism tests in line 5 can be done using Igusa invariants of a genus-2 curve (cf. [25]).
@@
The structure of the graph $\mathcal{G}_{2}(2,p)$ was studied by many authors (e.g. [28], [17]). Importantly, it is known (cf. [16, Theorem 7.2]) that its Jacobians’ subgraph $\mathcal{J}_{2}(2,p)$ is also connected. Here $\mathcal{J}_{g}(\ell,p)$ is defined to be
  • The vertices are isomorphism classes of Jacobians of superspecial curves of genus $g$ over $\overline{\mathbb{F}}_{p}$.
@@
The above mentioned research may enable us to construct a more efficient algorithm for listing superspecial genus-2 curves rather than Algorithm 2.21.
Similarly to the case of $g=2$, we want to list superspecial curves of genus $g\geq 3$. However, no explicit formula is known for the computations of $(2,\ldots,2)$-isogenies in general, unlike the genus-2 case. In the next section, we will propose an algorithm for computing $(2,2,2)$-isogenies using theta functions.
@@ 3. Computing the Richelot isogeny graph of dimension 3
In this section, we give an explicit algorithm (Algorithm 3.16) for listing superspecial genus-3 curves. For this purpose, we describe the following subroutines in the next three sections:
  • Section 3.1: For a given theta null-point of an abelian threefold $A$, compute a theta null-point of one $A^{\prime}$ which is $(2,2,2)$-isogenous to $A$.
@@
These computations use the theta constants introduced in Section 2.3, but they are still valid over a field of odd characteristic $p\geq 3$, thanks to the algebraic theory by Mumford [37]. In Section 3.4, we show that our main algorithm can be performed over $\mathbb{F}_{p^{4}}$, and provide a proof of Theorem 1.1 finally.
@@ 3.1. How to compute one path from the vertex
In this subsection, let $A$ be a complex abelian threefold, and $\varOmega\in\mathcal{H}_{3}$ be a period matrix of $A$. First, we shall prepare the following fact about a theta null-point of $A$.
@@
For a matrix $\varOmega\in\mathcal{H}_{3}$, the equation
$\vartheta_{0}\vartheta_{1}\vartheta_{2}\vartheta_{3}-\vartheta_{4}\vartheta_{5}\vartheta_{6}\vartheta_{7}=\vartheta_{32}\vartheta_{33}\vartheta_{34}\vartheta_{35}\vspace{0.5mm}$
holds with the notation in Definition 2.11.
@@
$m_{3}$ $\coloneqq\begin{pmatrix}0&0&0\\ 0&1/2&0\end{pmatrix},\quad$ $m_{4}$ $\coloneqq\begin{pmatrix}0&0&0\\ 1/2&1/2&0\end{pmatrix},\vspace{0.3mm}$
@@
$(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=\left\{\begin{array}[]{l}(1,1,1,0,0,0)\quad\text{if }\,k=1,\\[1.42262pt] (0,1,0,1,1,0)\quad\text{if }\,k=2,\\[1.42262pt] (0,1,0,0,0,0)\quad\text{if }\,k=3,\\[1.42262pt] (0,0,1,0,1,1)\quad\text{if }\,k=4,\\[1.42262pt] (0,0,1,0,0,0)\quad\text{if }\,k=5,\\[1.42262pt] (0,0,0,0,1,0)\quad\text{if }\,k=6\end{array}\right.$
@@
$\alpha\coloneqq\mathbf{1}_{3},\quad\beta\coloneqq\mathbf{0}_{3},\quad\gamma\coloneqq\begin{pmatrix}4c_{11}&2c_{12}&2c_{13}\\ 2c_{12}&4c_{22}&2c_{23}\\ 2c_{13}&2c_{23}&4c_{33}\end{pmatrix},\quad\delta\coloneqq\mathbf{1}_{3}.\vspace{0.5mm}$
It is obvious that $\alpha\hskip 0.56905pt^{t}\beta={\mathbf{0}}_{3},\,\gamma\hskip 0.85358pt^{t}\delta=\gamma$ are both symmetric and $\alpha\hskip 0.85358pt^{t}\delta-\beta\hskip 1.42262pt^{t}\gamma={\mathbf{1}}_{3}$ holds, hence $M$ belongs to ${\rm Sp}_{6}(\mathbb{Z})$ by Lemma 2.2. Moreover, one can check that $M$ is the desired matrix since
@@
(3.7) $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,2\varOmega)^{2}=\frac{1}{2^{3}}\hskip 0.56905pt\sum_{\beta\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)\theta\bigl{[}\begin{smallmatrix}0\\ b+\beta\end{smallmatrix}\bigr{]}(0,\varOmega)\vspace{0.7mm}$
for all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}$. Since the right-hand side of (3.7) is obtained by the values $\vartheta_{0},\ldots,\vartheta_{7}$, one can compute a theta null-point of $B$. As a summary of the statements in this subsection, we give an explicit algorithm (Algorithm 3.3) to compute a null-point of the codomain of a $(2,2,2)$-isogeny:
@@
It suffices to do the loops in lines 8–14 only for $i\in I_{\rm even}$, since we have that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$ for all $i\in I_{\rm odd}$, in fact.
@@ 3.2. How to compute all paths from the vertex
$\bigl{\{}M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{6}(\mathbb{Z})\mid\gamma\equiv\mathbf{0}_{3}\!\pmod{n}\bigr{\}}\vspace{0.3mm}$
@@
For a symplectic matrix $M\in{\rm Sp}_{6}(\mathbb{Z})$, it is well-known (cf. [7, §$\S 2.3.2$]) that $G$ is invariant (i.e. $M.G=G$) under the action by $M$ if and only if $M\in\varGamma_{0}(2)$. Hence, we need to act an element of ${\rm Sp}_{6}(\mathbb{Z})/\varGamma_{0}(2)$ on the theta null-point of $A$ in order to obtain a different $(2,2,2)$-isogeny from $A$.
Remark 3.6.
@@
In the following, let $\mathcal{M}$ be the output of the function CollectSymplecticMatrices. Then, we can compute theta null-points of $135$ abelian threefolds $(2,2,2)$-isogenous to $A$ by using Theorem 2.9 as in the following algorithm:
@@ 3.3. How to restore a defining equation of a genus-3 curve
  • If $N_{\rm van}=1$ and $S_{\rm van}\neq\{61\}$, how should we restore a defining equation of the corresponding genus-3 curve?
@@
$P_{i}\coloneqq\begin{pmatrix}\mathbf{1}_{3}&\beta_{i}\\[-0.85358pt] \gamma_{i}&\mathbf{1}_{3}\end{pmatrix},\vspace{-0.8mm}$
@@
    $\beta_{i}\coloneqq{\rm diag}(b_{1},b_{2},b_{3}),\quad\gamma_{i}\coloneqq{\rm diag}(a_{1},a_{2},a_{3}).\vspace{0.6mm}$
@@
[…]
@@
    $\beta_{62}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix}},\quad\gamma_{62}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}}.\vspace{0.3mm}$
@@
In the following, let $A$ be a principally polarized abelian threefold isomorphic to the Jacobian of a genus-3 hyperelliptic curve $C$ and suppose that a theta null-point of $A$ is given. By Proposition 2.15, there uniquely exists $i\in I_{\rm even}$ such that $\vartheta_{i}=0$. If $i=61$, we can apply Proposition 2.16 for restoring a defining equation of $C$, but otherwise not. Here, let $M:=P_{61}{P_{i}}^{-1}\in{\rm Sp}_{6}(\mathbb{Z})$ and $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ be the theta null-point after acting $M$ on given $[\vartheta_{i}^{2}]_{i}$, and we have ${\vartheta^{\prime}}_{\!\!61}=0$ by Lemma 3.11. Then one can apply Proposition 2.16 to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$, which enables us to recover $C$ as follows:
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=1$ of a p.p.a.v. $A$ of $\dim=3$
@@
Next, we suppose that a theta null-points of $A={\rm Jac}(C)$ is given, where $C$ is a smooth plane quartic. In the following, we explain how to recover $C$ according to the method in [35, §$\S 6.4$]. It follows from Theorem 2.6 that
@@
By using these equations, we can restore a defining equation of a plane quartic $C$, as follows:
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=0$ of a p.p.a.v. $A$ of $\dim=3$
@@
3.4. Algorithm for listing superspecial genus-3 curves
Let $p>2$ be a prime integer, and all principally polarized abelian varieties considered in this subsection are defined over a field of positive characteristic $p$. We emphasize that, even in this situation, all the algorithms until the previous subsection still work, thanks to the Mumford’s theory [37] (see also [11, §$\S 2$]). First, we prepare the following theorem:
Theorem 3.14.
Let $\hskip 0.56905ptC$ be a superspecial genus-$g$ curve in odd characteristic $p>2$. Then, the theta null-point of its Jacobian is $\mathbb{F}_{p^{2}}$-rational.